OpenStudy (ledah):

Can someone solve this as an example for me please? x+3y=0 and 20x-15y=75 using elimination.

OpenStudy (ajprincess):

Multiply the first eqn by 5 then u will get 5x+15y=0 nw add this with the second eqn. Then u will get 25x=75 x=75/25 =3. Substituting this value in the first eqn givs u 5+3y=0 3y=-5 y=-5/3 Is this clear or u want more explanation?

OpenStudy (ledah):

Thats clear enough thanks

OpenStudy (ajprincess):

yw